OpenStudy (anonymous):

the length of a rectangle is 7cm greater then its width.Find the length of the rectangle if its area is 120cm squared

OpenStudy (anonymous):

help

OpenStudy (gorv):

area=length*width

OpenStudy (gorv):

let width=x

OpenStudy (gorv):

length is 7 cm greater than width

OpenStudy (gorv):

length=x+7

OpenStudy (gorv):

120=x*(x+7)

OpenStudy (gorv):

x^2+7x=120 subtract 120 fromboth side

OpenStudy (gorv):

x^2+7x-120=0

OpenStudy (gorv):

x^2+15x-8x-120=0 now solve it further